OpenStudy (anonymous):

Solve 2x + y = 7 x – y = –4, please help, Its easy I just forgot exactly what to do

OpenStudy (anonymous):

solve for x and y either by using the elimination method, or gaussian elimination, or substitution, or etc

OpenStudy (blurbendy):

you can let x = y -4 then use substitution for x in the first equation

OpenStudy (anonymous):

do I need to multiply by the reciprocal to eliminate?

OpenStudy (blurbendy):

no, you can just add both equations, and the y's will cancel

OpenStudy (anonymous):

so 2x^2=3?

OpenStudy (anonymous):

okay thank you :)

OpenStudy (blurbendy):

2x + x = 3x

OpenStudy (blurbendy):

3x = 3 x = 1

OpenStudy (anonymous):

alright you just solved and got 3 then substituted right ?

OpenStudy (blurbendy):

x = 1, not 3, i made a mistake now you can substitute x = 1 in for the bottom equation 1 - y = -4 solve for y

OpenStudy (anonymous):

so 5 and 1

OpenStudy (anonymous):

-5

OpenStudy (blurbendy):

yes, x = 1 y = 5 you can put them back into either equation to check your work

OpenStudy (anonymous):

something is off wouldnt need to be 1-(-5) = -4 ?

OpenStudy (blurbendy):

no, y = 5 1 - (5) = - 4 -4 = -4

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

I got it now thanks a lot :D